what is the ratio of diffusion of the following:

A. Ne and Ar

B.PH3 and NH3

C.C4H10 and C2H4


1
Expert's answer
2022-05-06T03:30:02-0400

"\frac{r_1}{r_2}=\sqrt{\frac{M_2}{M_1}}"

A. "k=\sqrt{\frac{M(Ne)}{M(Ar)}}=\sqrt{\frac{20}{40}}=0.7"

B. "k=\sqrt{\frac{M(PH_3)}{M(NH_3)}}=\sqrt{\frac{31+3x1}{14+3x1}}=1.37"

C. "k=\sqrt{\frac{M(C_4H_{10})}{M(C_2H_4)}}=\sqrt{\frac{4x12+1x10}{2x12+4x1}}=1.44"